OpenStudy (anonymous):

Help please. Why is -1/e^(-1/2) = -e^(1/2)? (without using a calculator)

OpenStudy (anonymous):

A negative exponent means that something is "on the wrong side" of the fraction bar. \[x^{-1}=\frac{x^{-1}}{1}=\frac{1}{x}\] and \[\frac{1}{x^{-1}} = x\]

OpenStudy (anonymous):

Have to be careful with this, as \[\frac{1}{1+e^{-x}} \neq \frac{1+e^x}{1}\] or anything like that.

OpenStudy (anonymous):

but why did e^(-1/2) turned into -e^(1/2)

OpenStudy (anonymous):

It didn't. Not quite. You missed a negative sign on the left hand side.

OpenStudy (anonymous):

cause i saw this in cal book

OpenStudy (anonymous):

\[\frac{-1}{e^{\frac{-1}{2}}} = (-1)\frac{1}{e^{\frac{-1}{2}}} = (-1)(e^{\frac{1}{2}}) = -e^{\frac{1}{2}}\]

OpenStudy (anonymous):

oh...... lol that simple thx